Sketch the graph of the following function in one rectangular coordinate plane. Compare the behavior of each of the following graphs to the graph of y = sin x in terms of period, amplitude, and phase shift.




y=2sin x-4



y=sin 2x+3π2



y=sin 2x+2




1
Expert's answer
2022-02-28T15:52:00-0500

Explanations & answer


  • Here is the picture of the graphs attached. You can refer to it.



  • Y = Sin(x) is shown in the middle in dotted line.
  • Whenever there is a variation to that function it behaves like follows.
  • Y = 2Sin(x), function gets stretched along the Y-axis by a factor of 2 and Y = 2Sin(x) - 4 gets its middle axis shifted to the negative Y side (down) in 4 units. Hence the amplitude is increased by a factor of 2, with no phase shift and no change in period.


  • Y = sin(2x) gets its period reduced by a factor of 2. Then Y = Sin(2x+3"\small \pi"/2) gets a negative phase shift (shifts to left of x-axis). This would be the other way around if there was -3"\small \pi"/2 instead. Finally, there is no change in amplitude as the coefficient is 1 same as Y =Sin(x).



  • The last one is a combination of the previous two scenarios. It has its period reduced by 2 and the middle axis shifted in positive 2 along Y-axis (upwards). There is amplitude change as well as phase shift.